# Robust Optimization-based Autonomous Dynamic Soaring with a Fixed-Wing UAV

### v1

@@ Abstract
Dynamic soaring is a flying technique to exploit the energy available in wind shear layers, enabling potentially unlimited flight without the need for internal energy sources. We propose a framework for autonomous dynamic soaring with a fixed-wing unmanned aerial vehicle. The framework makes use of an explicit representation of the wind field and a classical approach for guidance and control of the unmanned aerial vehicle. Robustness to wind field estimation error is achieved by constructing point-wise robust reference paths for dynamic soaring and the development of a robust path following controller for the fixed-wing unmanned aerial vehicle. The framework is evaluated in dynamic soaring scenarios in simulation and real flight tests. In simulation, we demonstrate robust dynamic soaring flight subject to varied wind conditions, estimation errors and disturbances. Critical components of the framework, including energy predictions and path-following robustness, are further validated in real flights to assure small sim-to-real gap. Together, our results strongly indicate the ability of the proposed framework to achieve autonomous dynamic soaring flight in wind shear.
@@ I INTRODUCTION
Some bird species, such as the Wandering albatross can traverse over $5000\text{\,}\mathrm{km}$ in five days relying on wind energy alone [sachs2019maximum]. Heart rate monitoring and direct observation suggest that they use dynamic soaring to extract energy from the wind shear that occur over the ocean [pennycuick_1982, weimerskirch_2000]. Dynamic soaring and static soaring are the two primary methods to extract energy from the atmosphere. Static soaring utilizes rising air (often thermal updrafts) to gain energy with altitude, dynamic soaring extracts energy by traversing wind shear layers, i.e. vertical variations of the wind speed.
Extracting energy from the atmosphere would, analogously, provide a virtually unlimited range for an unmanned aerial vehicle, enabling extended-range missions and long-endurance persistent monitoring. Specifically, autonomous dynamic soaring with unmanned aerial vehicles has been gathering interest, as an alternative to thermal soaring as wind shears may be more widely available in the atmosphere than localized updrafts. Unlike thermal soaring, which has been demonstrated in multiple works [allen_guidance_2008, tabor_ardusoar_2018, depenbusch_autosoar_2018], however, full autonomous dynamic soaring has not yet been demonstrated on a real unmanned aerial vehicle. Although there is a broad body of research on the modeling and optimization of dynamic soaring trajectories [zhao2004optimal], only few lines of research consider actual demonstration for dynamic soaring in realistic wind fields. Works such as [bird2014closing, bronz_flight_2021] have demonstrated deployments of wind estimation and following an optimized path. However, to the best of our knowledge, energy-positive, sustained autonomous dynamic soaring has yet to be demonstrated.
@@ II RELATED WORK
Understanding how some bird species extract energy from the atmosphere has long been of interest [Rayleigh]. Optimization approaches have been used to understand the dynamic soaring process by finding the optimal trajectory to exploit energy from wind shear [zhao2004optimal, flanzer2012robust]. [bousquet_optimal_2017] showed that the optimal dynamic soaring consists of large shallow arcs, which are consistent with observations. [sachs_minimum_2005] estimates the minimum wind shear strength required for an Albatross to use dynamic soaring.
Recent research has demonstrated interest in improving unmanned aerial vehicle flight endurance by acquiring energy using thermal soaring [depenbusch_autosoar_2018, tabor_ardusoar_2018, notter2023deep] or solar power [AtlantikSolar]. In contrast, utilizing dynamic soaring for unmanned aerial vehicles has been relatively under-explored. [lawrance2011path] solves a planning problem to find an optimal path using discrete maneuver libraries, where the wind field is updated using Gaussian process regression. Reinforcement learning approaches have also been used to generate a policy for optimizing energy gain through dynamic soaring [montella_reinforcement_2014, chung2015learning, kim2019deep]. While there have been earlier attempts of dynamic soaring [bird2014closing, bronz_flight_2021], these works do not demonstrate dynamic soaring on a real platform in realistic scenarios.
One of the main challenges in enabling dynamic soaring on a real vehicle is estimating the highly uncertain wind field. Estimating a global, time-averaged wind field is challenging because it primarily depends on direct, in-situ measurements, with wind gusts and sensor noise compounding the problem. In [suys_autonomous_2023], orographic soaring in a wind tunnel is demonstrated. [bronz_simultaneous_2023] evaluates onboard wind estimates by comparing a ground-based Doppler lidar against wind estimates acquired by a quadrotor and an onboard multi-hole probe-based wind estimate. [mckenna_online_2023] considers the uncertainty of the wind estimates by incorporating the estimation uncertainty online for the trajectory optimization problem. [lawrance2011path] use Gaussian processes to model the uncertainty of the wind field.
In this work, we use a robust trajectory optimization approach to ensure that the optimized trajectory is robust to wind estimation errors. For tracking, we use an incremental nonlinear dynamic inversion similar to [Tal], where we show that the system is approximately differentially flat.
@@ III PROBLEM FORMULATION
Figure 2: A simplified dynamic soaring cycle can be described by four phases: upward wind climb (1) and downwind descent (3) are energy gain phases, while high altitude turn (2) and low altitude turn (4) describe energy loss phases.
Dynamic soaring is a technique to extract energy from wind shear, which are present in boundary layers close to the ground, or in leeward regions of geographic features blocking the wind. By choosing the path to repeatedly cross the shear layer in a direction that increases the airspeed, it is possible to compensate for aerodynamic drag or even perform energy-gain cycles. There are multiple types of DS cycles, e.g. wave-like or circular (Fig. 2). The main problem is to find a DS cycle that can reliably extract energy from the wind shear.
@@ III-A Wind Field Model
where $\mathbf{w}_{0}=[w_{0,x},w_{0,y},0]^{T}$, $z$ is the altitude, and $h$ and $s$ are parameters determining the vertical location and thickness of the shear layer, respectively. The sigmoid distribution $\sigma(z)$ has been adopted in previous work [bousquet_optimal_2017] to model a general shear layer.
@@ III-B Aircraft Model
Figure 3: Frame conventions used during this work. $\mathcal{I}$ denotes the inertial east north up frame, $\mathcal{B}$ denotes the body-fixed front right down frame and $\bm{\mathbf{v}}$ is the inertial velocity.
For the modeling of the aircraft dynamics, a simple lumped sum model is employed for the aerodynamics in combination with the Newton-Euler equations for a rigid body. The aerodynamic force $\bm{\mathbf{f}}$ acting on the aircraft can be modeled as Eq. (2), where $\rho$ is the air density, $\bm{\mathbf{v}}_{\text{air}}$} is the air-relative velocity, $A_{\text{wing}}$ is the planform wing area, $C_{L},C_{D}$ are normalized lift and drag coefficients, and $\alpha$ is the air-relative angle of attack of the wing [FlightMechanics].
@@
Similar to the aerodynamic model employed in [Tal], the aerodynamic moment contribution due to control surface deflections is modeled as in Eq. (5),
@@ IV DYNAMIC SOARING FRAMEWORK
The overview of the dynamic soaring framework is shown in Fig. 7. It consists of individual modules for the wind estimation, trajectory optimization and tracking control. The modules are structured such that uncertain wind fields are either explicitly accounted for a priori (robust path optimizer) or explicitly estimated (local wind estimator) and attenuated (path following controller).
@@ IV-A1 Local Wind Estimator
with $C_{A0}$, $C_{A1}$, $C_{B1}$ being aerodynamic coefficients. The aerodynamic forces $_{\mathcal{B}}\bm{\mathbf{f}}_{y,z}=\bm{\mathbf{a}}_{IMU,y,z}/m$ are derived from accelerometer readings. Further, $z_{\text{TAS}}$ is the true airspeed and $z_{\text{CAS}}$ is the calibrated airspeed. Finally, the wind velocity vector $\bm{\mathbf{w}}$ can be computed by applying a wind triangle computation with the estimated inertial velocity $\bm{\mathbf{v}}_{\text{ins}}$:
@@ IV-A2 Global Wind Estimator
The global wind estimator fuses local wind estimates to estimate the spatial distribution of the wind field. As the sigmoid function (1) is nonlinear in the parameters $\alpha$ and $h$, an extended Kalman filter makes use of local linearization of the observation dynamics. Therefore, the convergence guarantee of the linear Kalman Filter does not apply to this estimation problem [MPC]. Also, the wind field is only partially observable (e.g. an agent taking local wind field measurements must cover some subspace of the 3D space to estimate the wind field).
@@
$w_{x}$ $=w_{0,x}\cdot\sigma(z)+b_{x}\,\text{,}\quad w_{y}=w_{0,y}\cdot\sigma(z)+b_{y}$ (8)
$w_{z}$ $=0.$
The resulting state of the estimator is thus given by $\mathbf{x}=[w_{0,x},w_{0,y},b_{x},b_{y},s,h]^{T}$, and a trivial (random walk) process model is assumed for the wind field parameters due to the stationarity assumption. The estimate of the global wind field $w_{x},w_{y},s,h$ can then be used for trajectory optimization explained in the following subsection.
@@ IV-B Trajectory Optimization
We resort to a path-following approach, where the output of the trajectory optimizer is interpreted as a path. This spatial path, through different time parameterizations and states other than position, is mapped to a set of feasible trajectories. This allows to handle the infeasibility to track a trajectory due to the lack of thrust input. In the following formulation, the set of trajectories $\mathcal{X}$ corresponds to a set of different wind fields $\mathcal{W}$. Other sources of uncertainty, such as w.r.t. aerodynamic coefficients, can easily be accounted for in a similar way.
To setup the trajectory optimization, we employ the kinematic system model from [deittert2009engineless] with the state and input vector defined as
$\mathbf{x}$ $=[x,y,z,V,\gamma,\psi]^{T}$ (9)
$\mathbf{u}$ $=[C_{L},\mu]^{T}\text{,}$ (10)
@@
Figure 5: Robust energy gain paths, colored by the initial velocity at the lowest point of the path.
$J^{*}(\mathbf{x},$ $\mathbf{u},\mathbf{t})=\max_{\mathbf{x}_{0},\{\mathbf{u}_{i},\mathbf{t}_{i}\}_{i=0}^{N-1}}$ $\min_{\{j\}_{j=0}^{M}}V_{N}^{j}$ (11)
s.t. $\mathbf{x}_{i+1}^{j}=f_{j}(\mathbf{x}_{i}^{j},\mathbf{u}_{i}^{j},\mathbf{t}_{i}^{j})$ $\forall j$ $\in\{0,M\},\forall i\in\{0,N-1\}$
$\mathbf{x}_{i}^{j}\in\mathcal{C_{X}}$ $\forall j$ $\in\{0,M\},\forall i\in\{0,N\}$
$\mathbf{u}_{i}^{j}\in\mathcal{C_{U}}$ $\forall j$ $\in\{0,M\},\forall i\in\{0,N-1\}$
$\mathbf{x}_{0}^{j}\in\mathcal{C_{X}}_{0}$ $\forall j$ $\in\{0,M\}$
$\mathbf{x}_{N}^{j}\in\mathcal{C_{X}}_{N}$ $\forall j$ $\in\{0,M\}$
$S\mathbf{x}_{i}^{0}=S\mathbf{x}_{i}^{j}$ $\forall j$ $\in\{1,M\},\forall i\in\{0,N\}$
Note that $f_{j}(\cdot)$ denotes the system dynamics in the $j^{\text{th}}$ wind field. The common path constraint is enforced by the last constraint, with the selection matrix $S$. Note that we choose to maximize the worst-case terminal velocity $\forall\bm{\mathbf{w}}\in\mathcal{W}$. In practice, an equivalent but better-structured optimal control problem is used to pre-compute robust paths for a (uniform, grid) sampling of wind fields and initial velocities offline. Only the common reference path of the optimal solution of (11) is stored and then passed to the tracking controller. For differentiability of the path kinematics, we use a B-spline representation of the path with basis functions $\Phi=\{\phi_{0}(\tau),\phi_{1}(\tau),\ldots,\phi_{15}(\tau)\}$, where $\tau\in[0,1]$ is the normalized time variable during the dynamic soaring cycle. The basis is given by

### v2

# Towards Robust Optimization-Based Autonomous Dynamic Soaring with a Fixed-Wing UAV

@@ Abstract
Dynamic soaring is a flying technique to exploit the energy available in wind shear layers, enabling potentially unlimited flight without the need for internal energy sources. We propose a framework for autonomous dynamic soaring with a fixed-wing unmanned aerial vehicle. The framework makes use of an explicit representation of the wind field and a classical approach for guidance and control of the unmanned aerial vehicle. Robustness to wind field estimation error is achieved by constructing point-wise robust reference paths for dynamic soaring and the development of a robust path following controller for the fixed-wing unmanned aerial vehicle. Wind estimation and path tracking performance are validated with real flight tests to demonstrate robust path-following in real wind conditions. In simulation, we demonstrate robust dynamic soaring flight subject to varied wind conditions, estimation errors and disturbances. Together, our results strongly indicate the ability of the proposed framework to achieve autonomous dynamic soaring flight in wind shear.
@@ I INTRODUCTION
Some bird species, such as the Wandering albatross can traverse over $5000\text{\,}\mathrm{km}$ in five days relying on wind energy alone [18]. Heart rate monitoring and direct observation suggest that they use dynamic soaring to extract energy from the wind shear that occur over the ocean [16, 22]. Dynamic soaring and static soaring are the two primary methods to extract energy from the atmosphere. Static soaring utilizes rising air (often thermal updrafts) to gain energy with altitude, dynamic soaring extracts energy by traversing wind shear layers, i.e. vertical variations of the wind speed.
Extracting energy from the atmosphere would, analogously, provide a virtually unlimited range for an unmanned aerial vehicle, enabling extended-range missions and long-endurance persistent monitoring. Specifically, autonomous dynamic soaring with unmanned aerial vehicles has been gathering interest, as an alternative to thermal soaring as wind shears may be more widely available in the atmosphere than localized updrafts. Unlike thermal soaring, which has been demonstrated in multiple works [1, 8], however, full autonomous dynamic soaring has not yet been demonstrated on a real unmanned aerial vehicle. Although there is a broad body of research on the modeling and optimization of dynamic soaring trajectories [23], only few lines of research consider actual demonstration for dynamic soaring in realistic wind fields. Works such as [2, 4] have demonstrated deployments of wind estimation and following an optimized path. However, to the best of our knowledge, energy-positive, sustained autonomous dynamic soaring has yet to be demonstrated.
@@ II RELATED WORK
Understanding how some bird species extract energy from the atmosphere has long been of interest [17]. Optimization approaches have been used to understand the dynamic soaring process by finding the optimal trajectory to exploit energy from wind shear [23, 9]. [3] showed that the optimal dynamic soaring consists of large shallow arcs, which are consistent with observations. [19] estimates the minimum wind shear strength required for an Albatross to use dynamic soaring.
Recent research has demonstrated interest in improving unmanned aerial vehicle flight endurance by acquiring energy using thermal soaring [8, 14] or solar power [15]. In contrast, utilizing dynamic soaring for unmanned aerial vehicles has been relatively under-explored. [11] solves a planning problem to find an optimal path using discrete maneuver libraries, where the wind field is updated using Gaussian process regression. Reinforcement learning approaches have also been used to generate a policy for optimizing energy gain through dynamic soaring [13, 6]. While there have been earlier attempts of dynamic soaring [2, 4], these works do not demonstrate dynamic soaring on a real platform in realistic scenarios.
One of the main challenges in enabling dynamic soaring on a real vehicle is estimating the highly uncertain wind field. Estimating a global, time-averaged wind field is challenging because it primarily depends on direct, in-situ measurements, with wind gusts and sensor noise compounding the problem. In [20], orographic soaring in a wind tunnel is demonstrated. [5] evaluates onboard wind estimates by comparing a ground-based Doppler lidar against wind estimates acquired by a quadrotor and an onboard multi-hole probe-based wind estimate. [12] considers the uncertainty of the wind estimates by incorporating the estimation uncertainty online for the trajectory optimization problem. [11] use Gaussian processes to model the uncertainty of the wind field.
In this work, we take a different path by employing a robust strategy in both planning and control. A robust trajectory optimizer and an incremental nonlinear dynamic inversion tracking controller similar to [21] to ensure that the planned trajectories remain dynamically feasible, provide energy gain and can be tracked accurately in spite of significant, unknown perturbations in the wind field.
@@ III PROBLEM FORMULATION
Figure 2: A simplified dynamic soaring cycle is described by four phases: upward wind climb (1) and downwind descent (3) are energy gain phases, while high altitude turn (2) and low altitude turn (4) describe energy loss phases.
Dynamic soaring is a technique to extract energy from wind shear, which are present in boundary layers close to the ground, or in leeward regions of geographic features blocking the wind. There are multiple types of DS cycles, e.g. wave-like or circular (Fig. 2). The main problem is to find a DS cycle that can reliably extract energy from the wind shear.
@@ III-A Wind Field Model
where $\mathbf{w}_{0}=[w_{0,x},w_{0,y},0]^{T}$, $z$ is the altitude, and $h$ and $s$ are parameters determining the vertical location and thickness of the shear layer. The sigmoid distribution $\sigma(z)$ was adopted in previous work [3] to model a general shear layer.
@@ III-B Aircraft Model
Figure 3: Frame conventions, $\mathcal{I}$ denotes the inertial east north up frame, $\mathcal{B}$ denotes the body-fixed front right down frame, $\bm{\mathbf{v}}$ is the inertial velocity.
For the modeling of the aircraft dynamics, a simple lumped sum model is employed for the aerodynamics in combination with the Newton-Euler equations for a rigid body with body-fixed frame $\mathcal{B}$. The aerodynamic force $\bm{\mathbf{f}}$ acting on the aircraft can be modeled as Eq. (2), where $\rho$ is the air density, $\bm{\mathbf{v}}_{\text{air}}$} is the air-relative velocity, $A_{\text{wing}}$ is the planform wing area, $C_{L},C_{D}$ are normalized lift and drag coefficients, and $\alpha$ is the air-relative angle of attack of the wing [10].
@@
Similar to the aerodynamic model employed in [21], the aerodynamic moment contribution due to control surface deflections is modeled as in Eq. (5),
@@ IV DYNAMIC SOARING FRAMEWORK
The proposed dynamic soaring framework consists of individual modules for the wind estimation, trajectory optimization and tracking control (Fig. 4). The modules are structured such that uncertain wind fields are either explicitly accounted for a priori (robust path optimizer) or explicitly estimated (local wind estimator) and attenuated (path following controller).
@@ IV-A1 Local Wind Estimator
with $C_{A0}$, $C_{A1}$, $C_{B1}$ being aerodynamic coefficients , where the coefficients were determined from a nonlinear least-squares fit from flight data and are shown in Table II. The aerodynamic forces $_{\mathcal{B}}\bm{\mathbf{f}}_{y,z}=\bm{\mathbf{a}}_{IMU,y,z}/m$ are derived from accelerometer readings. Further, $z_{\text{TAS}}$ is the true airspeed and $z_{\text{CAS}}$ is the calibrated airspeed. Finally, the wind velocity vector $\bm{\mathbf{w}}$ can be computed by applying a wind triangle computation with the estimated inertial velocity $\bm{\mathbf{v}}_{\text{ins}}$:
@@ IV-A2 Global Wind Estimator
The global wind estimator fuses local wind estimates to estimate the spatial distribution of the wind field. As the sigmoid function (1) is nonlinear in the parameters $\alpha$ and $h$, an extended Kalman filter makes use of local linearization of the observation dynamics. Therefore, the convergence guarantee of the linear Kalman Filter does not apply to this estimation problem. Also, the wind field is only partially observable (e.g. an agent taking local wind field measurements must cover some subspace of the 3D space to estimate the wind field).
@@
$w_{x}=w_{0,x}\cdot\sigma(z)+b_{x}\,\text{,}\quad w_{y}=w_{0,y}\cdot\sigma(z)+b_{y}\,\text{,}\quad w_{z}=0.$ (8)
The resulting state of the estimator is thus given by $\mathbf{x}=[w_{0,x},w_{0,y},b_{x},b_{y},s,h]^{T}$, and a trivial (identity plus random walk) process model is assumed for the wind field parameters due to the stationarity assumption. We use identity matrices multiplied with $1e^{-6}$ and $4$ for the process and measurement covariances respectively and run the extended Kalman filter at $10\text{\,}\mathrm{Hz}$. The estimate of the global wind field $w_{x},w_{y},s,h$ is then used for trajectory optimization explained in the following subsection.
@@ IV-B Trajectory Optimization
We resort to a path-following approach, where the output of the trajectory optimizer is interpreted as a path. This spatial path is mapped to a set of feasible trajectories, through different time parameterizations and states other than position. This allows to handle the infeasibility of trajectory tracking resulting from the lack of a thrust input. In the following formulation, the set of trajectories $\mathcal{X}$ corresponds to a set of different wind fields $\mathcal{W}$. Other uncertainties, such as aerodynamic coefficients, can be accounted for in a similar way.
To setup the trajectory optimization, we employ the kinematic system model from [7] with the state and input vector
$\mathbf{x}=[x,y,z,V,\gamma,\psi]^{T},\quad\mathbf{u}=[C_{L},\mu]^{T}\text{,}$ (9)
@@
$J^{*}(\mathbf{x},$ $\mathbf{u},\mathbf{t})=\max_{\begin{subarray}{c}\mathbf{x}_{0},\Delta t,\{\mathbf{p}_{i}\}_{i=0}^{N},\\ \{\{\mathbf{u}_{i}^{j},\mathbf{t}_{i}^{j}\}_{i=0}^{N-1}\}_{j=0}^{M}\end{subarray}}V_{\text{min}}$ (10a)
s.t. $V_{\text{min}}\leq\{V_{N}^{j}\}_{j=0}^{M},\quad\forall j\in\{0,M\}$ (10b)
$\mathbf{x}_{i+1}^{j}=f_{j}(\mathbf{x}_{i}^{j},\mathbf{u}_{i}^{j},\mathbf{t}_{i}^{j})$ (10c)
$\hskip 60.00009pt\forall j\in\{0,M\},i\in\{0,N-1\}$
$S\mathbf{x}_{i}^{j}=\mathbf{p}_{i},\quad\forall j\in\{0,M\},i\in\{0,N\}$ (10d)
$t_{i}^{0}=\Delta t,\quad\forall i\in\{0,N\}$ (10e)
$\mathbf{x}_{i}^{j}\in\mathcal{C_{X}},\quad\forall j\in\{0,M\},i\in\{0,N\}$ (10f)
$\mathbf{u}_{i}^{j}\in\mathcal{C_{U}},\quad\forall j\in\{0,M\},i\in\{0,N-1\}$ (10g)
$\mathbf{x}_{0}^{j}\in\mathcal{C_{X}}_{0},\quad\mathbf{x}_{N}^{j}\in\mathcal{C_{X}}_{N},\quad\forall j\in\{0,M\}$ (10h)
The above problem effectively maximizes the worst-case terminal velocity for all wind fields $\forall\bm{\mathbf{w}}\in\mathcal{W}$ (10b), while the path constraint (10d) with the selection matrix $S$ extracting position to ensure that all trajectories share a common path. In (10c), $f_{j}(\cdot)$ denotes the system dynamics in the $j^{\text{th}}$ wind field. The various constraint sets are specified in Table I.
Figure 6: Combined feasibility and sensitivity analysis of the robust paths. The robust paths considering 4 perturbed wind fields remain dynamically feasible for significantly larger perturbations in the encountered wind field.
TABLE I: Constraint sets used in the optimization
Set Quantity Definition / Bound
$\mathcal{C_{U}}$ Lift $C_{L}$ $[0,1.0]$
Bank $\mu$ $\pi,0]$
Roll Rate $|\Delta\mu|\leq 2.0\cdot\Delta t$
$\mathcal{C_{X}}$ Speed $v$ $[10,100]$
Altitude $z$ $50,50],\ z_{k}\geq z_{0}$
Heading $\psi$ $\psi_{k+1}\leq\psi_{k}$
Time step $\Delta t$ $[0.05,1.0]$
$\mathcal{C_{X}}_{0}$ Pos. $\mathbf{p}_{0}$ $0,0,z_{0}]^{T},\ z_{0}\in[-50,0]$
Angles $\gamma_{0},\psi_{0}$ 0
$\mathcal{C_{X}}_{N}$ State $\mathbf{p}_{N}=\mathbf{p}_{0},\ \gamma_{N}=\gamma_{0}$
Heading $\psi_{N}$ $\psi_{0}-2\pi$
We use the optimal control problem (10) to pre-compute robust paths for a (uniform, grid) sampling of nominal wind fields with 4 perturbed wind fields each ($M=5$). The increased robustness under uncertainty is shown in Fig. 6, where some efficiency is traded for robust feasibility of a fixed path under constraints (Table I) and uncertainty in the wind field. Note that this differs from the feasibility of the constrained, nominal problem under a known wind field. We use the IPOPT solver with implicit midpoint integration and multiple shooting in our implementation to compute the robust paths for varying initial velocities in an offline step. Only the common reference path of the optimal solution of (10) is stored and then passed to the tracking controller. For differentiability of the path kinematics, we use a B-spline representation of the path with basis functions $\Phi=\{\phi_{0}(\tau),\phi_{1}(\tau),\ldots,\phi_{15}(\tau)\}$, where $\tau\in[0,1]$ is the normalized time variable during the dynamic soaring cycle. The basis [^1] The chosen basis is $C^{\infty}$ continuous $\forall t\in(0,1)$, but only $C^{0}$ continuous at $t=0\rightarrow t=1$. A Bernstein basis is an alternative choice here. is given by